OpenStudy (anonymous):

Write an equation of a line perpendicular to line AB below in slope-intercept form that passes through the point (7, 6).

OpenStudy (anonymous):

y = 0.5x + 2.5 y = 0.5x - 2.5 y = -2x + 20 y = 2x - 20

OpenStudy (anonymous):